# AuthGlass: Enhancing Voice Authentication on Smart Glasses via Air-Bone Acoustic Features

### v2

@@ 4.1. Overview of the Authentication Structure
To begin, we define an authentication framework based on prior research(Chang, 2018; Liu et al., 2018; Huang et al., 2025). Since our primary focus is on analyzing the acoustic features captured by smart glasses, we adopt a simple structure as the fundamental design. Fig. LABEL:fig:authentication_structure illustrates our authentication model, which consists of three main modules: data pre-processing, liveness detection, and user authentication.
@@ 4.2.1. Principle for Embedding Algorithm
After primary features were extracted, they were then embedded to a user-related feature either for enrollment or authentication, as described in section 4.1. Previous works have proposed different methods for embedding and comparison during authentication, such as Principal Component Analysis(Iwano et al., 2005), Gaussian mixture model(Liu et al., 2018), and Deep Neural Networks(Huang et al., 2025). As we aim to evaluate the effectiveness of various features, the embedding algorithm should be compatible with different input modalities. Furthermore, the algorithm needs to be independent of feature type, treating all modalities uniformly. Therefore, deep neural network (DNN) methods are the most suitable choice, owing to their flexibility and strong capability to extract embeddings for distinguishing different inputs.
@@ 4.3.1. Method
We performed a 7-fold cross-user evaluation on the data collected from 42 subjects. In each fold, the data of 36 subjects were used for training with a 10-times data augmentation on the audio amplitude, while the remaining 6 subjects were reserved for testing authentication accuracy. For enrollment, each test user provided one sample per command to simulate real-world usage, and the system averaged the corresponding embeddings to construct that user’s template. Then, we evaluated the performance of authentication using Equal Error Rate (EER) (Huang et al., 2025)as metric, as it gives a balanced measurement on both the false acceptance and false reject. Given the limited dataset size and our focus on exploring the potential of different features, no validation set was employed. Each training trial ran for 50 epochs, and the authentication EER was evaluated every 5 epochs. Model that reached lowest EER during 50 epochs were selected and the EER metrics as long as the threshold for cosine-similarity calculation were averaged across 7 folds.
@@ 4.3.2. Result and Discussion
We calculated the averaged EER and threshold of different feature combination across 7 folds of experiments. which achieved at the lowest EER as 0.03 with a 0.62 threshold. This demonstrate the effectiveness of our method.
@@ 5.1. Attack Model
Similar with section 4, we need to firstly define the attack model. With the rapid advancement of audio generative methods(Kamble et al., 2020; Wenger et al., 2021), solely rely on audio features to defend against certain kinds of adversarial sample could not guarantee the effectiveness of the authentication system. Therefore, we designed the attack evaluation based on two assumptions. (a) the attacker can perfectly capture and replay the air-conductive signals and (b) the attacker could be able to simulate the speech production process, but cannot perfectly replicate that of each individuals. A total of three attack experiments were conducted.
@@ 5.2. Experiment 1: Attack by Simulated Wearer
Considering when an attacker could access the smart glasses device, we aimed to use an simulator to attack the authentication system. Thus, as shown in Fig 3a, we incorporated GRAS-45BC-KEMAR(KEM, 2025), a torso-mouth simulator to generate adversarial samples. The simulator could not only produce high-quality audio but could also simulate the exhaling process and torso vibration during speaking, making it a perfect example to generate audio samples as well as the vibration and sound field features.
@@
Then, with these adversarial samples, we attack the models trained in section 4.3 separately following the 7-fold group. For each groups of data, we followed the enrollment process same as those in 4.3 via one original sample per command, and then attack using the collected attack samples divided into folds respectively. We utilized Attack Success Rate (ASR) to evaluate the system’s robustness against attack. Result shows that our method could effectively defend wearer attack.
@@ 5.3. Experiment 2: Attack by Nearby Sound Source
Apart from direct attack by simulated wearer, voice authentication on smart glasses could suffer from attack from sound source from environment, as the devices are exposed to surrounding voices. Therefore, we utilized PreSonus Eris E5(Eri, 2025), a high-quality speaker as the attacker sound source. To control the relative position between user and attacker, we leveraged the simulator used in previous experiment as the user with fixed position when putting the smart glasses on. Then, we selected six position located on four direction of the simulated user, as shown in Fig 3b. Similar with experiment 1, we utilized all the genuine samples for attack and balanced the volume for different attacking points. After getting the result following the 7-fold division. Since all the PR achieved to 0, we only averaged the ASR of all attacking position for evaluation. Result also shows that our method is effective in defending attack.
@@ 5.4. Experiment 3: Synthetic Attack Using Original Audios
Although we incorporated high quality method to generate attack samples and tested the defending capability of the proposed authentication system in experiment 1 and 2, there’s still a chance that the adversarial samples did not accurately reconstruct the original audio features. Thus, we substituted the AC features in all attack samples of previous experiments with the genuine ones, and utilized these synthesized features as attack samples and re-evaluate the performance of the authentication system. By conducting the evaluation, we could gain further understanding of whether vibration signals and sound-field features could enhance the authentication performance, as the adversarial samples share the same AC features with the genuine data. For both the wearer and the nearby sound source scenario, our method demonstrate a relative ability to resist attacks originating from surrounding voices.
@@ 6. Usability of AuthGlass
Figure 4. In the restricted microphone experiment. (a) shows the three microphones selected. (b) shows the five microphones selected.(c) presents the authentication accuracy using different number of microphones. Yellow indicates inputs containing only AC microphones, while blue indicates inputs containing both AC and BC microphones.
@@ 6.1. Restricted Microphone Number
Aiming to exploit a throughout acoustic features by smart glasses, we embedded 14 AC microphones and 2 BC ones, which were redundant compared with commercial smart glasses. Realizing this, we investigated whether leveraging a limited number of microphones could achieve a competitive authentication performance. Based on various designs of modern smart glasses, we selected two representative cases: a 3-channel microphone array such as Rokid(Rok, 2025) and a 5-channel microphone array such as Rayban Meta(Met, 2025). To evaluate the performance of these two kinds of configurations, we selected the specific AC microphone combination on our prototype to simulate the microphone layout, as shown in Fig 4. In these cases, the microphone number was restricted and sound field feature could only be partially captured. Then, we calculated the EER among 7-folds of data to evaluate the performance under limited microphone settings. As shown in Fig 4. Result shows that as the channel number increased, the performance using sound field features increase accordingly, while BC signals have better performance compared with sound field features. Also, three channel sound features did not yielded better performance due to the insufficient sound field information.
@@ 6.2. Cross-Session Validation
In practical applications on smart glasses, authentication systems are expected to deliver stable results across different wearing sessions. Unlike controlled laboratory settings, users may wear the device at varying times or under slightly different conditions. Thus, to ensure reliability, the system must exhibit cross-session robustness, meaning that the authentication outcome remains consistent every time the device is worn. To validate the cross-session performance, we recruited 23 of the 42 subjects in section 3.2.2 again, and asked them to redo the same process as that during data-collection. Then, using the model that has already trained, we conducted the evaluation steps similar in section 4.3, leveraging the old data for enrollment (1 sample per command) while using the newly collected data to authenticate. The results demonstrate that compared with the original EER obtained in the initial session, there is virtually no performance degradation.
@@ 6.3.1. Method
For deployment on smart glasses, authentication systems inevitably face various sources of noise. Thus, to evaluate the robustness under noisy conditions, we introduced environmental noise into the input signals. Since our prototype captures data from synchronized multiple-channel AC microphones placed at certain layout, it is not appropriate to incorporate noises dataset from mono microphone. Therefore, we collected real-world ambient sounds from three typical scenarios: cafés, restaurants, and shopping malls using our smartglasses prototype, while simultaneously recording the sound levels with a decibel meter. After collecting the noises, we mixed the noise with origin signals at certain degrees of signal-to-noise ratio (SNR), using channel 7 microphone as reference. With the data mixed with noises, we further tested the authentication performance by EER metric using the same model as 4.3 following the 7-fold settings. Result shows that when the SNR decreases, the performance of all methods is degraded to a very high EER indicating that the system could be affected by high noises.

### v3

@@ 4. Using Air-Bone Features for Liveness Authentication
Section 3.3 provides a thorough analysis of potential features for smart-glasses liveness authentication. However, their practical applicability and effectiveness remain unclear. We therefore instantiate a simple, representative authentication pipeline and empirically evaluate these features.
Figure 7. Pre-processing exploits the features and utilizes a feature embedding algorithm to generate a data template. During authentication, features are passed to the liveness detection module and authentication algorithm compare the embedded data to the template.
@@ 4.1. Overview of the Authentication Structure
To begin, we define an authentication framework based on prior research(chang2018two; Vocal_resonance; huang2025eve). Since our primary focus is on analyzing the acoustic features captured by smart glasses, we adopt a simple structure as the fundamental design. Fig. 7 illustrates our authentication model, which consists of three main modules: data pre-processing, liveness detection, and user authentication.
@@
4.2. Selected Features During Pre-processing
To evaluate the effectiveness of different features, we selected representative features for AC signal (voiceprint), vibration, and sound field, repectively. Firstly, we chose channel 7 (Fig 2a) AC microphone as baseline, since it captured most detailed air-conducted audio signal. To extract the voiceprint of this microphone, we incorporate common-used log Mel-spectrogram (meghanani2021exploration; nguyen2023heart) as the representative feature. Then, as demonstrated in section 3.3, vibration signals captured by BC microphones showed consistency with AC signals. Thus, we incorporate log Mel-spectrogram to process BC data, similar with AC feature. Finally, we opted TDM and EDM feature for sound field, as it speaks to the energy change and time delay variance during sound propogation.
4.3. Liveness Detection Through Vibration and Sound Field Features
Although prior studies have shown that a single AC microphone can be used for liveness detection(ahmed2020void; robustDetect), their methods largely rely on artifacts specific to artificial speakers and may break down when facing different speaker types or adversarial strategies(qu2022synthesising). By contrast, leveraging speech-related biometric features offers greater robustness, as machines cannot fully replicate the human speech production process. Thus, based on findings from section 3.3, we utilized both the consistency between AC and BC signals, and a uniformed pattern of TDM and EDM across all users for liveness detection.
4.3.1. Consistency between AC and BC signal
Figure 8. Each box in the boxplot represents the interquartile range from the 0.25 to the 0.75 quantile, the red line indicates the median, the yellow dots denote outliers, and the red dashed line marks the 0.01 quantile. For (a) cosine similarity of Mel-spectrograms between AC and BC speech segments, the 0.01 quantile is 0.985, while for (b) the consistency score based on method from Huang et al(huang2025eve) with parameter M = 4, N = 4, the 0.01 quantile is 0.942.
To evaluate the consistency between AC and BC signal, we computed the correlation between the log Mel-spectrogram of these two signals at the same time. Fig. 8a shows the correlation scores across all sentences and users, where most cases exhibit high values, demonstrating a strong consistency of vibration and air-conducted acoustic signals when user is speaking. As Huang et al(huang2025eve) proposed an alternative algorithm regarding the consistency, we leveraged this method for comparison with the algorithm parameters configured as M = 4 and N = 4, as shown in Fig 8. To balance the accuracy of the liveness detection system and its capability of rejecting adversarial samples, we set the acceptance rate at 99% and implemented the threshold for detection.
4.3.2. Uniformed TDM and EDM pattern
Figure 9. Each box in the boxplot represents the interquartile range from the 0.25 to the 0.75 quantile, the red line indicates the median, the yellow dots denote outliers, and the red dashed line marks the 0.01 quantile. For (a) TDM cosine similarity between TDM and uniformed template, the 0.01 quantile for all sample is 0.991, while for (b) EDM cosine similarity, the 0.01 quantile is 0.978.
Since we have revealed the energy change and time delay variance along spatial dimension in section 3.3, we then tried to further incorporate it for liveness detection, as it speaks to where the received signal is from. To achieve this, we averaged the TDM and EDM along the time axis of each channels, forming a mean TDM and EDM template (similar with Fig 5b). Then we calculated the mean template across all users and all samples, resulting an uniformed TDM and EDM pattern. Finally, we calculated the cosine similarity between the uniformed pattern and template from all subjects, as shown in Fig 9. As all users exhibited high correlation value, we set the acceptance rate at 99% as the threshold with same consideration as section before.
@@ 4.4.1. Principle for Embedding Algorithm
After primary features were extracted, they were then embedded to a user-related feature either for enrollment or authentication, as described in section 4.1. Previous works have proposed different methods for embedding and comparison during authentication, such as Principal Component Analysis(iwano2005multimodal), Gaussian mixture model(Vocal_resonance), and Deep Neural Networks(huang2025eve). As we aim to evaluate the effectiveness of various features, the embedding algorithm should be compatible with different input modalities. Furthermore, the algorithm needs to be independent of feature type, treating all modalities uniformly. Therefore, deep neural network (DNN) methods are the most suitable choice, owing to their flexibility and strong capability to extract embeddings for distinguishing different inputs.
Figure 10. The embedding model consists of four parts: head, ResNet backbone, command head and speaker head. The head allow 1-3 input features and we adjust the first layer of ResNet convolution to fit the input. Command head employs a gradient reversal layer to obfuscate command distinctions, making the model usable across all commands. Speaker head output the authentication result.
4.4.2. Embedding Algorithm Design
During usage, users are required to authenticate using different voice commands, such as ’Pay the Bill’ and ’Call Tom’. Therefore, it is necessary to employ a DNN model capable of accommodating various sentence inputs. Previous studies (sun2024eyegesener; huang2025eve) have shown that models based on ResNet (he2016deep) combined with domain-adversarial learning (DAL) (tzeng2017adversarial; ganin2015unsupervised) achieve satisfactory accuracy with unsupervised domain adaption, which could be used for authentication across different commands. Based on this, as shown in Fig 10, we incorporated all convolution layers in ResNet18 as backbone structure and projected the output to a 256-dimension embedding by a fully connected layer. The embedding was then passed to two heads: the speaker head and the command head, representing the detected speaker and the command, respectively. To adapt to different combination of features, each input features processed by respective convolution head were concatenated together and then passed to the backbone model, where the input channel was expanded accordingly.
The training process of the embedding model served as a classification problem. During training, the model was fed with labeled data from the training set and optimized by minimizing the combined entropy loss from the speaker and command heads. The latter employs a gradient reversal layer to obfuscate command distinctions, thereby encouraging the DNN to learn embeddings that remain usable across all commands. After training, we discarded the speaker and command heads and retained only the embedding layer to generate user embeddings for enrollment and authentication, assuming that the shared backbone captures representative information of individual characteristics regardless of the specific command.
@@ 4.5.1. Method
After designing the liveness authentication process, we aimed to investigate whether the selected features could support authentication and to assess their relative performance. To this end, we defined several feature groups, using the case that relies solely on the AC feature as the baseline, since it directly reflects the speaker’s voiceprint. We then assessed the contribution of vibration and sound field features when combined with AC signals, denoted as AC + BC and AC + DMs (where DMs refer to TDM + EDM), respectively. Finally, we evaluated the performance using all three types of features—AC, vibration, and sound field—denoted as AC + BC + DMs. In addition, as a training baseline, we used the straightforward sound field feature—the log Mel-spectrograms from all channels except the reference AC channel—denoted as SF, and conducted experiments similar to those with DMs.
For each feature group, we performed a 7-fold cross-user evaluation on the data collected from 42 subjects. In each fold, the data of 36 subjects were used for training with a 10-times data augmentation on the audio amplitude, while the remaining 6 subjects were reserved for testing authentication accuracy. For enrollment, each test user provided one sample per command to simulate real-world usage, and the system averaged the corresponding embeddings to construct that user’s template. Then, we evaluated the performance of authentication using Equal Error Rate (EER) (huang2025eve)as metric, as it gives a balanced measurement on both the false acceptance and false reject. Given the limited dataset size and our focus on exploring the potential of different features, no validation set was employed. Each training trial ran for 50 epochs, and the authentication EER was evaluated every 5 epochs. Model that reached lowest EER during 50 epochs were selected and the EER metrics as long as the threshold for cosine-similarity calculation were averaged across 7 folds.
AC AC+DMs AC+SF AC+BC AC+BC+DMs AC+BC+SF
mean EER 0.09 0.05 0.08 0.05 0.03 0.06
mean threshold 0.49 0.61 0.53 0.58 0.62 0.66
Table 1. Averaged EER and threshold of Different Feature Combination.
@@ 4.5.2. Result and Discussion
Table 1 shows the averaged EER and threshold of different feature combination across 7 folds of experiments. As illustrated, AC + BC and AC + DMs combination yields approximate 0.05 decrease on EER compared with single AC group, indicating that both vibration signal (BC) and sound field features could be leveraged to enhance authentication performance respectively. Furthermore, combining the BC and DMs feature could achieve better result with even lower EER. In contrast, using SF features as an end-to-end methods proved to be not that effective, as AC + SF group only has a slight EER decrease compared with single AC ones and there is no difference between AC + BC + SF group and the AC + BC ones. Thus, we leverage DMs as the sound field feature as it can capture both energy changes and time delay variance.
@@ 5.1. Attack Model
Similar with section 4, we need to firstly define the attack model. With the rapid advancement of audio generative methods(kamble2020advances; wenger2021hello), solely rely on audio features to defend against certain kinds of adversarial sample could not guarantee the effectiveness of the authentication system. Therefore, we designed the attack evaluation based on two assumptions. (a) the attacker can perfectly capture and replay the air-conductive signals and (b) the attacker could be able to simulate the speech production process, but cannot perfectly replicate that of each individuals. A total of three attack experiments were conducted.
@@ 5.2. Experiment 1: Attack by Simulated Wearer
5.2.1. Method
Considering when an attacker could access the smart glasses device, we aimed to use an simulator to attack the authentication system. Thus, as shown in Fig 11a, we incorporated GRAS-45BC-KEMAR(KEMAR), a torso-mouth simulator to generate adversarial samples. The simulator could not only produce high-quality audio but could also simulate the exhaling process and torso vibration during speaking, making it a perfect example to generate audio samples as well as the vibration and sound field features.
@@ 5.2.1. Method
Then, with these adversarial samples, we attack the models trained in section 4.5 separately following the 7-fold group. For each groups of data, we followed the enrollment process same as those in 4.5 via one original sample per command, and then attack using the collected attack samples divided into folds respectively. To evaluate the effectiveness of the methods for liveness detection in section 4.3 we calculated the pass rate (PR) of the attack samples by these methods. For simplicity, we denoted the consistency between AC and BC signals calculated by cosine similarity as CS1, while the method from Huang et al(huang2025eve) as CS2. For the method using uniformed TDM and EDM pattern, we just denoted them as TDM and EDM respectively. Though there was only partial of attack samples passed through liveness detection, we still fed all attack samples to the authentication system for comprehensive evaluation. We utilized Attack Success Rate (ASR) to evaluate the system’s robustness against attack.
CS1 CS2 TDM EDM
PR 0.004 0.998 0 0.998
(a) PR for four liveness detection metrics
AC AC+BC AC+DMs AC+BC+DMs
ASR 0.1596 0.0077 0.0491 0.0003
(b) ASR for different feature combinations
Table 2. Averaged Passing Rate (PR) and Attack Success Rate (ASR) for experiment 1.
5.2.2. Result and Discussion
Table 2 demonstrated the PR for four types of liveness methods and the ASR for four types of feature combination, following the configurations in section 4.5. Since SF feature did not yield significant promotion on authentication performance, we did not incorporated this feature for ASR calculation. As shown in the table, CS1 yields lower pass rate, meaning that it’s suitable for smart glass scenario compared with CS2, which only gives rough estimation on the consistency of AC and BC signals.
As shown in table 2b, both BC signals and sound field feature (DM) promote the resilience against attack, while combining them together reached the lowest ASR rate. This indicating that vibration signals and sound field features could both be contributive for authentication.
@@ 5.3. Experiment 2: Attack by Nearby Sound Source
5.3.1. Method
Apart from direct attack by simulated wearer, voice authentication on smart glasses could suffer from attack from sound source from environment, as the devices are exposed to surrounding voices. Therefore, we utilized PreSonus Eris E5(Eris), a high-quality speaker as the attacker sound source. To control the relative position between user and attacker, we leveraged the simulator used in previous experiment as the user with fixed position when putting the smart glasses on. Then, we selected six position located on four direction of the simulated user, as shown in Fig 11b. Similar with experiment 1, we utilized all the genuine samples for attack and balanced the volume for different attacking points. After getting the result following the 7-fold division. Since all the PR achieved to 0, we only averaged the ASR of all attacking position for evaluation.
AC AC+BC AC+DMs AC+BC+DMs
ASR 0.0960 0.0310 0.0338 0.0204
Table 3. Attack Success Rate (ASR) for experiment 2.
5.3.2. Result and Discussion
Table 3 demonstrated the PR for four types of liveness detection and ASR for four types of feature combination, following the configurations in section4.5. For liveness detection, both the AC-BC consistency reached a relatively low performance with 0.65 passing rate, we concluded that the bone-conductive microphone were affected by the air-conducted signal. Since the audio collection room incorporated noise canceling configurations, the AC channel could be consistent with the BC signal. For TDM and EDM template, TDM yielded best performance as the signal is not from the mouth, leading to a total different TDM pattern. While for EDM, it cannot capture Similar with previous experiment, combining DM and BC yielded the best performance.
@@ 5.4. Experiment 3: Synthetic Attack Using Original Audios
5.4.1. Method
Although we incorporated high quality method to generate attack samples and tested the defending capability of the proposed authentication system in experiment 1 and 2, there’s still a chance that the adversarial samples did not accurately reconstruct the original audio features. Thus, we substituted the AC features in all attack samples of previous experiments with the genuine ones, and utilized these synthesized features as attack samples and re-evaluate the performance of the authentication system. By conducting the evaluation, we could gain further understanding of whether vibration signals and sound-field features could enhance the authentication performance, as the adversarial samples share the same AC features with the genuine data. Since the evaluations on liveness detection were conducted before, we only considered the ASR of different combination of features, and neglected the AC group since we leveraged the actual AC feature for authentication.
AC+BC AC+DMs AC+BC+DMs
Simulated Wearer 0.0046 0.1046 0.0167
Nearby Sound Source 0.0306 0.0408 0.0027
Table 4. Averaged Attack Success Rate (ASR) for experiment 3.
5.4.2. Result and Discussion
Table 4 demonstrated the ASR for three types of feature combination. As shown in the table, incorporating genuine AC features substantially increases the ASR for AC+DMs in the simulated-wearer case, indicating that DM features alone are not sufficiently discriminative to block attacks when the simulated wearer closely resembles the genuine user. In contrast, for the nearby sound source scenario, both DMs and BC demonstrate a relative ability to resist attacks originating from surrounding voices.
@@ 6.1. Restricted Microphone Number
Aiming to exploit a throughout acoustic features by smart glasses, we embedded 14 AC microphones and 2 BC ones, which were redundant compared with commercial smart glasses. Realizing this, we investigated whether leveraging a limited number of microphones could achieve a competitive authentication performance. Based on various designs of modern smart glasses, we selected two representative cases: a 3-channel microphone array such as Rokid(RokidGlasses2025) and a 5-channel microphone array such as Rayban Meta(MetaRayban2025). To evaluate the performance of these two kinds of configurations, we selected the specific AC microphone combination on our prototype to simulate the microphone layout, as shown in Fig 12. In these cases, the microphone number was restricted and sound field feature could only be partially captured. Despite this, considering the effectiveness of EDM and TDM feature in representing sound field, we still leveraged the corresponding channels of AC microphone to compute these two features. Then, we retraced the evaluation steps in section 4.5, including embedding model training, enrollment and authentication process. Then, we calculated the EER among 7-folds of data to evaluate the performance under limited microphone settings. As shown in Fig 12. Result shows that as the channel number increased, the performance using sound field features increase accordingly, while BC signals have better performance compared with sound field features. Also, three channel sound features did not yielded better performance due to the insufficient sound field information.
@@ 6.2. Cross-Session Validation
6.2.1. Method
In practical applications on smart glasses, authentication systems are expected to deliver stable results across different wearing sessions. Unlike controlled laboratory settings, users may wear the device at varying times or under slightly different conditions. Thus, to ensure reliability, the system must exhibit cross-session robustness, meaning that the authentication outcome remains consistent every time the device is worn. To validate the cross-session performance, we recruited 23 of the 42 subjects in section 3.2.2 again, and asked them to redo the same process as that during data-collection. Then, using the model that has already trained, we conducted the evaluation steps similar in section 4.5, leveraging the old data for enrollment (1 sample per command) while using the newly collected data to authenticate.
6.2.2. Result and Discussion
Table 5 demonstrated the Averaged EER of Different Feature Combination for Cross-Session Validation. The results demonstrate the averaged EER for different feature combinations in the cross-session validation. As shown in the table, all four feature sets achieve relatively high accuracy. Compared with the original EER obtained in the initial session, there is virtually no performance degradation, indicating that the authentication model maintains strong cross-session robustness and can consistently deliver reliable results across different wearing instances.
AC AC+DMs AC+BC AC+BC+DMs
mean EER 0.0981 0.0449 0.0861 0.0505
Table 5. Averaged EER of Different Feature Combination for Cross-Session Validation.
Figure 12. In the restricted microphone experiment, TDM and EDM were calculated using three and five microphones, respectively. (a) shows the three microphones selected. (b) shows the five microphones selected.(c) presents the authentication accuracy using different number of microphones. Yellow indicates inputs containing only AC microphones, while blue indicates inputs containing both AC and BC microphones.
@@ 6.3.1. Method
For deployment on smart glasses, authentication systems inevitably face various sources of noise. Thus, to evaluate the robustness under noisy conditions, we introduced environmental noise into the input signals. Since our prototype captures data from synchronized multiple-channel AC microphones placed at certain layout, it is not appropriate to incorporate noises dataset from mono microphone. Therefore, we collected real-world ambient sounds from three typical scenarios: cafés, restaurants, and shopping malls using our smartglasses prototype, while simultaneously recording the sound levels with a decibel meter. After collecting the noises, we mixed the noise with origin signals at certain degrees of signal-to-noise ratio (SNR), using channel 7 microphone as reference. With the data mixed with noises, we further tested the authentication performance by EER metric using the same model as 4.5 following the 7-fold settings.
6.3.2. Result and Discussion
As shown in Table 6, when the SNR decreases, the performance of all methods is degraded to a very high EER. Instead, for 10 SNR group, combining BC signals or sound field features did not enhance the performance, as the sound field features and BC audios could be affected when there are noises. Finally, for 20 SNR group, combining all the features yielded better performance than other groups, indicating that the system could be affected by high noises.
SNR (dB) AC AC+DMs AC+BC AC+BC+DMs
5 0.1491 0.1843 0.1593 0.1781
10 0.1046 0.1528 0.1417 0.1546
20 0.0898 0.0747 0.0849 0.0481
Table 6. Averaged EER of Different Feature Combination for Noisy Environment under Different SNR Levels.